there are 3 resistors in parallel with resistances of 3Ω,4Ω and 11Ω. What is the total resistance of the parallel combination?
1
Expert's answer
2021-10-04T19:38:36-0400

"\frac{1}{R}=\frac{1}{R_1}+\frac{1}{R_2}+\frac{1}{R_3}"

"\frac{1}{R}=\frac{1}{3}+\frac{1}{4}+\frac{1}{11}"

"R=\frac{44+33+12}{12\times11}=\frac{89}{132}=0.6742\Omega"